OpenStudy (anonymous):

A linear function L is defined so that L(5)=7 & L(-1)=25. Find a formula for L of the form L(x)=mx+b. Then graph L, showing both intercepts & labeling them with their coordinates.

OpenStudy (nenadmatematika):

L-7=18/(-6)*(x-5) so L=-3x+22